OpenStudy (anonymous):

VERIFY THE IDENTITY: 1/csc(x-1) + 1/csc(x+1)=2/csc(x)-sin(x)

OpenStudy (freckles):

question is it really: \[\frac{1}{\csc(x-1)}+\frac{1}{\csc(x+1)}=\frac{2}{\csc(x)}-\sin(x)\]

OpenStudy (anonymous):

close, only it says 2/csc(x)-sin(x) not sure if that is the same thing or not?

OpenStudy (freckles):

isn't that what I typed for the right hand side?

OpenStudy (freckles):

did you mean to write 2/(csc(x)-sin(x)) instead?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[\frac{1}{\csc(x-1)}+\frac{1}{\csc(x+1)}=\frac{2}{\csc(x)-\sin(x)}\]

OpenStudy (anonymous):

yes that is correct

OpenStudy (freckles):

that doesn't seem like an identity

OpenStudy (freckles):

try inputting pi/6

OpenStudy (freckles):

both sides aren't the same

OpenStudy (anonymous):

Ok that is what I kept coming up with, I just wanted to verify that it was not an identity, thank you!

OpenStudy (freckles):

did you mean to write \[\frac{1}{\csc(x)-1}+\frac{1}{\csc(x)+1}=\frac{2}{\csc(x)-\sin(x)}\] this is an identity

OpenStudy (anonymous):

I cannot figure out how it is an identity, however....

OpenStudy (freckles):

so is that not what you are trying to prove is an identity ? it really is the first equation you wrote?

OpenStudy (freckles):

If you really meant to write that last thing I wrote, try combining the fractions on the left hand side first

OpenStudy (anonymous):

I think I am typing it wrong, this is exactly how it is written on the worksheet: \[\frac{ 1 }{ \csc x-1 }+\frac{ 1 }{ \csc x+1 }=\frac{ 2 }{ \csc x-\sin x }\]

OpenStudy (freckles):

then all you have to do is divide both top and bottom by csc(x)

OpenStudy (freckles):

2 steps really: combine fractions divide top and bot by csc(x)

OpenStudy (anonymous):

How can you combine the fractions when the denominator are not the same? I am not sure how to make them the same

OpenStudy (unklerhaukus):

\[LHS=\frac{1}{\csc(x)-1}+\frac{1}{\csc(x)+1}\\ =\frac{1}{\csc(x)-1}\times\frac{\csc(x)+1}{\csc(x)+1}+\frac{1}{\csc(x)+1}\times\frac{\csc(x)-1}{\csc(x)-1}\\ =\frac{2\csc(x)}{\csc^2(x)-1}\\ =\] now divide both numerator and denominator by csc

OpenStudy (anonymous):

Like this? \[\frac{ 2 }{ \csc x-1}\]

OpenStudy (unklerhaukus):

\[\frac{ 2 }{ \csc x-\tfrac1{\csc(x)}}\]

OpenStudy (anonymous):

ohh of course. I understand now. Thank you so much!

OpenStudy (unklerhaukus):

all clear? do you understand how we combined those fractions at the start?

OpenStudy (anonymous):

yes, I kept trying to do it that way but for some reason I was doing the algebra wrong....I need to practice more